# Water wave scattering by a surface-mounted rectangular anisotropic elastic plate

### v2

@@ Abstract
This paper considers the problem of water wave scattering by a rectangular anisotropic elastic plate mounted on the ocean surface, with either free or clamped edges. The problem is obtained as an expansion over the dry modes of the elastic plate, which are computed using a Rayleigh–Ritz method. In turn, the component diffraction and radiation problems are solved by formulating a boundary integral equation and solving numerically using a constant panel method. The results are presented to highlight the resonant responses of the plate under different forcing scenarios. In particular, we illustrate how the excitation of certain modes can be forbidden due to symmetry.
@@ 1 Introduction
Hydroelasticity, which is concerned with the interaction of fluids with elastic bodies, is an important topic with numerous applications, including those in the cryosphere (ocean wave scattering by sea ice and ice shelves) and those related to the engineering of very large floating structures (VLFS). A comprehensive account of the historical research can be found in the review articles squire2007ocean; squire2020ocean. We give here a brief account of some of the hydroelastic models. We note that this extensive research has focussed almost entirely on isotropic plates.
@@
The two-dimensional water wave problem (one horizontal and one vertical) was solved by an eigenfunction expansion method by fox_squire94 and extended by sahoo2001scattering who analysed a semi-infinite elastic plate floating on the surface of water of finite-depth. The Wiener-Hopf technique is particularly suited to semifinite problems and highly analytic solutions have been derived in this case (balmforth1999ocean; chung2002calculation; tkacheva2003plane; williams_meylan12; smith2020wiener). The finite elastic plate was solved by meylan_squire94 using Green’s function technique, which led to a Fredholm integral equation and was simultaneously solved using a modal expansion by newman1994wave. The three-dimensional water wave problem involving circular elastic discs was solved by jgrfloecirc; peter_meylan_chung04. The solution for plates of arbitrary shapes was given in AppliedOR01; meylan2002wave. Similar methods were developed by hermans2004interaction; andrianov2006hydroelastic.
@@
Recently, extensive studies on hydroelasticity have been driven by its potential applications in wave energy conversion, with flexible wave energy converters (WECs) continuing to be a focal point of current research (Babarit2017; Philen2018; Tay2020; Collins2021; Renzi2021; Michele2022; Zheng2022; Vipin_Koley_2022; Guo2023; singh2023performance; Cheng2024; MEYLAN2025120931). A promising mechanism for flexible wave energy conversion is the piezoelectric effect, in which materials become electrically polarised in response to elastic deformation. First comprehensively modelled by renzi2016hydroelectromechanical, a piezoelectric WEC (usually modelled as a plate) would become polarised in response to deformation by ocean waves and thus induce a current in an external circuit. An important aspect of piezoelectric materials that has been hitherto ignored in the piezoelectric WEC literature is their anisotropy—while piezoceramics are usually transversely isotropic (Yang_2018), the polymer PVDF, which is often proposed for piezoelectric WECs, are anisotropic (Vinogradov01041999). To date, authors have bypassed considerations of anisotropy by focussing on the one-dimensional WEC/two-dimensional fluid context, with submerged horizontal plates (renzi2016hydroelectromechanical; Vipin_Koley_2022), surface mounted plates (MEYLAN2025120931) and vertical plates (singh2023performance) being among the cases considered. The present paper seeks to investigate wave scattering by a two-dimensional anisotropic elastic plate in a three-dimensional fluid. In doing so, this paper will facilitate future extensions of earlier models of piezoelectric WECs (renzi2016hydroelectromechanical; MEYLAN2025120931) to three dimensions.
@@
The outline of this paper is as follows. In §2 we describe a model of a rectangular anisotropic elastic plate mounted on the water surface. Our method is an extension of that presented by MEYLAN2025120931 for the two-dimensional case, in which the solution is expanded as a sum of a diffraction potential and a superposition of radiation potentials. The diffraction potential, which assumes that the plate is fixed, is solved in §3. The modes of vibration of the plate and the corresponding radiation potentials are found in §4 and §5, respectively. The coupled fluid/plate problem is solved in §6. An energy balance identity used to verify our computations, which is based on the optical theorem, is derived in §7. A large number of results are presented in §8, which serve to illustrate the capabilities of the model and act as benchmark solutions for future studies. A conclusion is given in §9.
@@ 2.1 Preliminaries
The plate, which has uniform thickness $h$, is assumed to be thin, which means $a,b\gg h$. The motion of the plate is modelled using Kirchhoff–Love theory. In variational form, the motion is governed by
$\iint_{\Gamma}\rho h\,\delta W\,\partial_{t}^{2}W+\mathcal{B}(\delta W,W)-P\delta W\,\mathrm{d}x\,\mathrm{d}y=0,$ (4)
where $W(x,y,t)$ is the vertical displacement of the plate, $\delta W$ is an admissible variation of $W$, $\rho$ is the uniform density of the plate and $P(x,y,t)$ is the fluid pressure under the plate (grossi2003; Reddy_2006; grossi2008). The symmetric bilinear operator $\mathcal{B}$ is given by
$\mathcal{B}(\delta W,W)=\begin{bmatrix}\partial_{x}^{2}\delta W&\partial_{y}^{2}\delta W&\partial_{x}\partial_{y}\delta W\end{bmatrix}\begin{bmatrix}D_{11}&D_{12}&2D_{16}\\ D_{12}&D_{22}&2D_{26}\\ 2D_{16}&2D_{26}&4D_{66}\end{bmatrix}\begin{bmatrix}\partial_{x}^{2}W\\ \partial_{y}^{2}W\\ \partial_{x}\partial_{y}W\end{bmatrix}$, (5)
where $D_{ij}$ are the rigidity coefficients which describe the material’s resistance to bending and twisting (Reddy_2006). The strong form of (4) is
@@
Note that in the isotropic case, $D_{11}=D_{22}=D_{12}+2D_{66}>0$ and $D_{16}=D_{26}=0$, in which case $\mathcal{D}$ is proportional to the biharmonic operator.
@@
for $(x,y)\in\Gamma$. By equating the fluid velocity at the fluid-plate interface with the plate velocity, this condition assumes that the plate remains fully in contact with the fluid at all times. We note that linear hydroelasticity is well validated experimentally for sufficiently low ratios of amplitude to wavelength (Montiel2013a; Montiel2013b).
@@
where $\Lambda=\{(0,0),(a,0),(0,b),(a,b)\}$ are the corners of the plate (grossi2003). We note that while our framework can be extended to other boundary condition types (e.g. simply supported edges, moored edges) or to have different boundary conditions on different edges, we do not explore this here.
@@ 2.2 Time harmonic formulation
for $(x,y)\in\Gamma$, where the plate satisfies the clamped boundary conditions (11). In the Sommerfeld condition (14e), $r=\sqrt{x^{2}+y^{2}}$, $k$ is the positive real root of the dispersion relation
@@ 3.2 Integral equation for diffraction problem
We note that the same integral equation may be obtained by using the Green’s function for a submerged source (as given by Linton2001, Equation B.86) and applying Green’s third identity to $\phi^{\rm di}$. This involves a subtle point which arises when the source point in on the free surface and we replace the normal derivative of the Green function using the free-surface condition. This point is discussed in jgrfloecirc; meylan2002wave where a similar Green’s function method is used and also in buchner1993evaluation.
@@ 3.3 Numerical solution using a constant panel method
for all $i$ (meaning this quantity only needs to be calculated once), where the second line follows from translation. The double integral is then calculated by (i) directly integrating the term $1/r$ in (19), and (ii) using the MATLAB routine integral2 to numerically integrate the remaining part. We remark that in preliminary tests of our method, the integral representation of the Green’s function (19) converged faster than methods based on an equivalent series representation (i.e., one analogous to Equation B.91 in Linton2001), presumably due to direct integration of the $1/r$ term that dominates in the near field. Our numerical code for the diffraction problem was validated for the case of a circular plate, using a code developed to solve the problem of wave scattering by a partially submerged truncated circular cylinder, which is valid in the limit of negligible submergence (wilks2025simultaneouslayoutdeviceparameter). The aforementioned code uses separation of variables in cylindrical coordinates to solve the problem both beneath and exterior to the plate. Subsequently, continuity conditions between the interior and exterior regions reduce to integral equations that are solved numerically using a singularity respecting Galerkin method.
@@ 4 Dry modes of vibration of the plate
and the boundary conditions (either (11) or (12) for a free or clamped plate, respectively), where $\omega_{j}$ are the angular frequencies of the vibrational modes. They are called dry modes because they solve the dynamic equation of the plate (6) in the absence of any forcing from the fluid (i.e. with $P=0$). The corresponding weak form is
@@
where $\delta\mathrm{w}$ is an admissible variation of $\mathrm{w}_{j}$. We seek a numerical solution of (30) using the Rayleigh–Ritz method, which is classical but included here for completeness (see Reddy_2006, for further discussion). To begin, we introduce a set of functions $\{\psi_{1},\dots,\psi_{N_{RR}}\}$, which satisfy the boundary conditions (11). We approximate $w_{j}$ in the form of a superposition over this set, namely
@@
where $\kappa_{m}$ are the beam eigenvalues and $\delta_{mn}$ is the Kronecker delta. Further details about these eigenmodes are given in Appendix B. As considered by Young1950, the basis for the Rayleigh–Ritz method is then taken to be
@@
where $\theta$ indexes all pairs $(m,n)\in\mathbb{N}^{2}$ such that $1\leq m\leq N_{x}$ and $1\leq n\leq N_{y}$, in which $N_{x}$ and $N_{y}$ are the number of beam modes used in the $x$ and $y$ dimension, respectively. Thus $N_{RR}=N_{x}N_{y}$. We note that this choice of basis functions is advantageous since they are mutually orthogonal and satisfy the plate boundary conditions (11), although other choices of basis functions could be made (e.g. polynomials, see Reddy_2006, Chapter 7). Arguments of $\theta$ in indices $m$ and $n$ will be dropped in what follows.
@@
$\iint_{\Gamma}w_{m}(x,y)w_{n}(x,y)^{*}\,\mathrm{d}x\mathrm{d}y=\delta_{mn}.$ (37)
@@
We take $N_{x}/a=N_{y}/b=20$ throughout this paper. We have validated our method for computing the dry modes using results provided by LEISSA1973257 in the case where the plate is isotropic, and results provided in 10.1115/1.4053090 in the case where the plate is anisotropic, with excellent agreement—see the first and third columns of Figure 2.
@@
Table 1: Rigidity coeffficients used to generate the results in this paper. The Poisson’s ratio associated with the isotropic plate rigidity coefficients is $0.3$. The rigidity coefficients of the orthotropic plate correspond to a material under plane-stress conditions for which $E_{y}/E_{x}=3/4$, $G_{xy}/E_{x}=1/2$ (where $E_{x}$ and $E_{y}$ are the Young’s moduli in the $x$ and $y$ directions, respectively, and $G_{xy}$ is the shear modulus) and the Poisson’s ratio is $0.3$. The coefficients for the anisotropic plate are as in 10.1115/1.4053090. All units are in ${\rm Pa}\,{\rm m}^{3}$.

### v3

@@ Abstract
This paper considers the problem of water wave scattering by a rectangular anisotropic elastic plate mounted on the ocean surface, with either free, clamped or simply-supported edges. The problem is obtained as an expansion over the dry modes of the elastic plate, which are computed using a Rayleigh–Ritz method. In turn, the component diffraction and radiation problems are solved by formulating a boundary integral equation and solving numerically using a constant panel method. The results are presented to highlight the resonant responses of the plate under different forcing scenarios. In particular, we illustrate how the excitation of certain modes can be forbidden due to symmetry.
@@ 1 Introduction
Hydroelasticity, which is concerned with the interaction of fluids with elastic bodies, is an important topic with numerous applications, including those in the cryosphere (ocean wave scattering by sea ice and ice shelves) and those related to the engineering of very large floating structures (VLFS). A comprehensive account of the historical research can be found in the review articles Squire (2007, 2020). We give here a brief account of some of the hydroelastic models. We note that this extensive research has focussed almost entirely on isotropic plates, with a notable exception being the recent work of Thery et al. (2026), who used a space transformation approach to determine the parameters of an anisotropic plate mounted over shallow water for cloaking applications.
@@
The two-dimensional water wave problem (one horizontal and one vertical) was solved by an eigenfunction expansion method by Fox and Squire (1994) and extended by Sahoo et al. (2001) who analysed a semi-infinite elastic plate floating on the surface of water of finite-depth. The Wiener-Hopf technique is particularly suited to semifinite problems and highly analytic solutions have been derived in this case (Balmforth and Craster, 1999; Chung and Fox, 2002; Tkacheva, 2003; Williams and Meylan, 2012; Smith et al., 2020). The finite elastic plate was solved by Meylan and Squire (1994) using Green’s function technique, which led to a Fredholm integral equation and was simultaneously solved using a modal expansion by Newman (1994). The three-dimensional water wave problem involving circular elastic discs was solved by Meylan and Squire (1996); Peter et al. (2004). The solution for plates of arbitrary shapes was given in Meylan (2001, 2002). Similar methods were developed by Hermans (2004); Andrianov and Hermans (2006).
@@
Recently, extensive studies on hydroelasticity have been driven by its potential applications in wave energy conversion, with flexible wave energy converters (WECs) continuing to be a focal point of current research (Babarit et al., 2017; Philen et al., 2018; Tay and Wei, 2020; Collins et al., 2021; Renzi et al., 2021; Michele et al., 2022; Zheng et al., 2022; Valappil and Koley, 2022; Guo et al., 2023; Singh and Gayen, 2023; Cheng et al., 2024; Meylan et al., 2025). A promising mechanism for flexible wave energy conversion is the piezoelectric effect, in which materials become electrically polarised in response to elastic deformation. First comprehensively modelled by Renzi (2016), a piezoelectric WEC (usually modelled as a plate) would become polarised in response to deformation by ocean waves and thus induce a current in an external circuit. An important aspect of piezoelectric materials that has been hitherto ignored in the piezoelectric WEC literature is their anisotropy—while piezoceramics are usually transversely isotropic (Yang, 2018), the polymer PVDF, which is often proposed for piezoelectric WECs, are anisotropic (Vinogradov and Holloway, 1999). To date, authors have bypassed considerations of anisotropy by focussing on the one-dimensional WEC/two-dimensional fluid context, with submerged horizontal plates (Renzi, 2016; Valappil and Koley, 2022), surface mounted plates (Meylan et al., 2025) and vertical plates (Singh and Gayen, 2023) being among the cases considered. The present paper seeks to investigate wave scattering by a two-dimensional anisotropic elastic plate in a three-dimensional fluid. In doing so, this paper will facilitate future extensions of earlier models of piezoelectric WECs (Renzi, 2016; Meylan et al., 2025) to three dimensions.
@@
The outline of this paper is as follows. In §2 we describe a model of a rectangular anisotropic elastic plate mounted on the water surface. Our method is an extension of that presented by Meylan et al. (2025) for the two-dimensional case, in which the solution is expanded as a sum of a diffraction potential and a superposition of radiation potentials. The diffraction potential, which assumes that the plate is fixed, is solved in §3. The modes of vibration of the plate and the corresponding radiation potentials are found in §4 and §5, respectively. The coupled fluid/plate problem is solved in §6. An energy balance identity used to verify our computations, which is based on the optical theorem, is derived in §7. A large number of results are presented in §8, which serve to illustrate the capabilities of the model and act as benchmark solutions for future studies. A conclusion is given in §9.
@@ 2.1 Preliminaries
The plate, which has uniform thickness $h$, is assumed to be thin, which means $a,b\gg h$. The motion of the plate is modelled using Kirchhoff–Love theory. We assume that in-plane deformations can be neglected because these will be negligible compared to bending. The Lagrangian of the plate is
$\mathcal{L}\{W\}=T\{W\}-U_{\rm{strain}}\{W\}-U_{\rm{pressure}}\{W\}$ (4)
where $W(x,y,t)$ is the vertical displacement of the plate. The kinetic energy, strain energy due to bending and potential energy due to pressure are
$T\{W\}$ $=\frac{1}{2}\rho h\iint_{\Gamma}(\partial_{t}W)^{2}\mathrm{d}x\,\mathrm{d}y$ (5a)
$U_{\rm{strain}}\{W\}$ $=\frac{1}{2}\iint_{\Gamma}\mathcal{B}(W,W)\mathrm{d}x\,\mathrm{d}y$ (5b)
$U_{\rm{pressure}}\{W\}$ $=-\iint_{\Gamma}PW\mathrm{d}x\,\mathrm{d}y,$ (5c)
respectively, $\rho$ is the uniform density of the plate and $P(x,y,t)$ is the fluid pressure under the plate (Grossi and Nallim, 2003; Reddy, 2006; Grossi and Nallim, 2008). The symmetric bilinear operator $\mathcal{B}$ is given by
$\mathcal{B}(W_{1},W_{2})=\begin{bmatrix}\partial_{x}^{2}W_{1}&\partial_{y}^{2}W_{1}&\partial_{x}\partial_{y}W_{1}\end{bmatrix}\begin{bmatrix}D_{11}&D_{12}&2D_{16}\\ D_{12}&D_{22}&2D_{26}\\ 2D_{16}&2D_{26}&4D_{66}\end{bmatrix}\begin{bmatrix}\partial_{x}^{2}W_{2}\\ \partial_{y}^{2}W_{2}\\ \partial_{x}\partial_{y}W_{2}\end{bmatrix}$, (6)
where $D_{ij}$ are the rigidity coefficients which describe the material’s resistance to bending and twisting (Reddy, 2006). The action over an arbitrary time interval $(t_{0},t_{1})$ and its first variation are
$\mathcal{A}$ $=\int_{t_{0}}^{t_{1}}\mathcal{L}\{W\}\mathrm{d}t$ (7)
$\delta\mathcal{A}$ $=\int_{t_{0}}^{t_{1}}\delta\mathcal{L}\{\delta W,W\}\mathrm{d}t$ (8)
$=\int_{t_{0}}^{t_{1}}\int_{\Gamma}\left\{\rho h\,\delta W\,\partial_{t}^{2}W+\mathcal{B}(\delta W,W)-P\delta W\right\}\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}t,$ (9)
respectively, where $\delta W$ is an admissible variation of $W$. The principle of least action implies that $\delta\mathcal{A}=0$ for arbitrary admissible variations $\delta W$ over arbitrary time intervals $(t_{0},t_{1})$, which yields
$\iint_{\Gamma}\left\{\rho h\,\delta W\,\partial_{t}^{2}W+\mathcal{B}(\delta W,W)-P\delta W\right\}\,\mathrm{d}x\,\mathrm{d}y=0.$ (10)
The strong form of (10) is
@@
Note that in the isotropic case, we have
$D_{11}=D_{22}$ $=D,$ (13a)
$D_{12}$ $=\nu D,$ (13b)
$D_{66}$ $=\tfrac{1}{2}(1-\nu)D,$ (13c)
$D_{16}=D_{26}$ $=0,$ (13d)
$\text{where }D$ $=\frac{Eh^{3}}{12(1-\nu^{2})},$ (13e)
in which $E$ is Poisson’s ratio and $\nu$ is the Young’s modulus. In this case, the operator $\mathcal{D}$ is proportional to the biharmonic operator.
@@
for $(x,y)\in\Gamma$. By equating the fluid velocity at the fluid-plate interface with the plate velocity, this condition assumes that the plate remains fully in contact with the fluid at all times. We note that linear hydroelasticity is well validated experimentally for sufficiently low ratios of amplitude to wavelength (Montiel et al., 2013a, b).
@@
where $\Lambda=\{(0,0),(a,0),(0,b),(a,b)\}$ are the corners of the plate (Grossi and Nallim, 2003), or simply supported, implying (§6.1.2 of Reddy, 2006)
$W=D_{11}\partial_{x}^{2}W+D_{12}\partial_{y}^{2}W$ $=0$ $x=0\text{ or }x=a,$ (19a)
$W=D_{12}\partial_{x}^{2}W+D_{22}\partial_{y}^{2}W$ $=0$ $y=0\text{ or }y=b.$ (19b)
We note that while our framework can be extended to other boundary condition types (e.g. moored edges) or to have different boundary conditions on different edges, we do not explore this here.
@@ 2.2 Time harmonic formulation
for $(x,y)\in\Gamma$, where the plate satisfies either clamped (17), free (18) or simply-supported (19) boundary conditions, as prescribed. In the Sommerfeld condition (21e), $r=\sqrt{x^{2}+y^{2}}$, $k$ is the positive real root of the dispersion relation
@@ 3.2 Integral equation for diffraction problem
We note that the same integral equation may be obtained by using the Green’s function for a submerged source (as given by Linton and McIver, 2001, Equation B.86) and applying Green’s third identity to $\phi^{\rm di}$. This involves a subtle point which arises when the source point in on the free surface and we replace the normal derivative of the Green function using the free-surface condition. This point is discussed in Meylan and Squire (1996); Meylan (2002) where a similar Green’s function method is used and also in Buchner (1993).
@@ 3.3 Numerical solution using a constant panel method
for all $i$ (meaning this quantity only needs to be calculated once), where the second line follows from translation. The double integral is then calculated by (i) directly integrating the term $1/r$ in (26), and (ii) using the MATLAB routine integral2 to numerically integrate the remaining part. We remark that in preliminary tests of our method, the integral representation of the Green’s function (26) converged faster than methods based on an equivalent series representation (i.e., one analogous to Equation B.91 in Linton and McIver (2001)), presumably due to direct integration of the $1/r$ term that dominates in the near field. Our numerical code for the diffraction problem was validated for the case of a circular plate, using a code developed to solve the problem of wave scattering by a partially submerged truncated circular cylinder, which is valid in the limit of negligible submergence (Wilks et al., 2025). The aforementioned code uses separation of variables in cylindrical coordinates to solve the problem both beneath and exterior to the plate. Subsequently, continuity conditions between the interior and exterior regions reduce to integral equations that are solved numerically using a singularity respecting Galerkin method.
@@ 4 Dry modes of vibration of the plate
and the boundary conditions (either (17), (18) or (19) for a clamped, free or simply-supported plate, respectively), where $\omega_{j}$ are the angular frequencies of the vibrational modes. They are called dry modes because they solve the dynamic equation of the plate (11) in the absence of any forcing from the fluid (i.e. with $P=0$). The corresponding weak form is
@@
where $\delta\mathrm{w}$ is an admissible variation of $\mathrm{w}_{j}$. We seek a numerical solution of (37) using the Rayleigh–Ritz method, which is classical but included here for completeness (see Reddy, 2006, for further discussion). To begin, we introduce a set of functions $\{\psi_{1},\dots,\psi_{N_{RR}}\}$, which satisfy the boundary conditions (17). We approximate $\mathrm{w}_{j}$ in the form of a superposition over this set, namely
@@
where $\kappa_{m}$ are the beam eigenvalues and $\delta_{mn}$ is the Kronecker delta. Further details about these eigenmodes are given in Appendix B. As considered by Young (1950), the basis for the Rayleigh–Ritz method is then taken to be
@@
where $\theta$ indexes all pairs $(m,n)\in\mathbb{N}^{2}$ such that $1\leq m\leq N_{x}$ and $1\leq n\leq N_{y}$, in which $N_{x}$ and $N_{y}$ are the number of beam modes used in the $x$ and $y$ dimension, respectively. Thus $N_{RR}=N_{x}N_{y}$. We note that this choice of basis functions is advantageous since they are mutually orthogonal and satisfy the plate boundary conditions (17), although other choices of basis functions could be made (e.g. polynomials, see Reddy, 2006, Chapter 7). Arguments of $\theta$ in indices $m$ and $n$ will be dropped in what follows.
@@
$\iint_{\Gamma}\mathrm{w}_{m}(x,y)\mathrm{w}_{n}(x,y)^{*}\,\mathrm{d}x\mathrm{d}y=\delta_{mn}.$ (44)
@@
Finally, in the case that the edges of the plate are simply supported, we construct the functions used in the Rayleigh–Ritz method from the modes of vibration of a simply supported-simply supported Euler–Bernoulli beam of unit length, which, when normalised, satisfy
$\partial_{x}^{4}u_{m}$ $=\kappa_{m}^{4}u_{m},$ (46a)
$u_{m}(0)$ $=u_{m}(1)=0,$ (46b)
$u_{m}^{(2)}(0)$ $=u_{m}^{(2)}(1)=0,$ (46c)
$\int_{0}^{1}u_{m}(x)u_{n}(x)\mathrm{d}x$ $=\delta_{mn}.$ (46d)
We take $N_{x}/a=N_{y}/b=20$ throughout this paper. We have validated our method for computing the dry modes using results provided by Leissa (1973, Table C1) in the case where the plate is isotropic, and results provided in An et al. (2021) in the case where the plate is anisotropic, with excellent agreement—see the first and third columns of Figure 2.
@@
Table 1: Rigidity coeffficients used to generate the results in this paper. The Poisson’s ratio associated with the isotropic plate rigidity coefficients is $0.3$. The rigidity coefficients of the orthotropic plate correspond to a material under plane-stress conditions for which $E_{y}/E_{x}=3/4$, $G_{xy}/E_{x}=1/2$ (where $E_{x}$ and $E_{y}$ are the Young’s moduli in the $x$ and $y$ directions, respectively, and $G_{xy}$ is the shear modulus) and the Poisson’s ratio is $0.3$. The coefficients for the anisotropic plate are as in An et al. (2021). All units are in ${\rm Pa}\,{\rm m}^{3}$.